OpenStudy (anonymous):

An astronaut on a strange planet finds that she can jump a maximum range of 15m with an initial speed of 3m/s. What is the free fall acceleration on the planet?

OpenStudy (anonymous):

0=3^2-2*a*15 9=30a a=9/30=.3m/s^2 that is the acceleration

OpenStudy (anonymous):

third law of motion: v^2=u^2-2as where a acceleration and s the distance u initial velocity and v final velocity